Solve the system by substitution.
x + y = 12
y = 2x
This problem has me a bit confused and need some help. Thanks.

Answer by stanbon(48568)

x + y = 12
y = 2x
----------
You are told that y=2x so substitute "2x" in place of "y"
in the 1st equation:
You get: x+2x=12
3x=12
x=4
Now you know what "x" is.
Substitute that value of x into the statement "y=2x"
You get y=2(4)=8
---------
So the solution for the system of equations is (4,8)
-------------
Cheers,
Stan H.